# and the homomorphism conjecture for $\mathcal{A}_{g}$

### v1

@@ Abstract.
was constructed in [9]. We present here new calculations of intersection products of the Torelli locus in $\mathcal{A}_{g}$ with the product loci
$\mathcal{A}_{r}\times\mathcal{A}_{g-r}\rightarrow\mathcal{A}_{g}$
for $r\leq 3$. The results suggest that $\mathsf{taut}$ is a $\mathbb{Q}$-algebra homomorphism (at least for special cycles). We discuss a conjectural framework for the homomorphism property of tautological projection.
@@
We define new tautological projections of cycles on the fiber products $\mathcal{X}_{g}^{s}\to\mathcal{A}_{g}$ of the universal family. We compute these projections for a class of product cycles on $\mathcal{X}_{g}^{s}$ in terms of a determinant involving the universal theta divisors and Poincaré classes. The Torelli pullback of the morphism $\mathcal{A}_{1}\times\mathcal{A}_{5}\to\mathcal{A}_{6}$ was used in [10] to construct a nontrivial element of the Gorenstein kernel of ${\mathsf{R}}^{5}(\mathcal{M}^{\mathrm{ct}}_{6})$. Using Abel-Jacobi pullbacks of product cycles on $\mathcal{X}_{g}^{s}$ and their projections, we construct a new family of classes which we conjecture to lie in the Gorenstein kernels of the tautological rings $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})$. In particular, we construct a nontrivial element of the Gorenstein kernel of $\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathrm{ct}})$. Whether these classes generate the Gorenstein kernels of all $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ is an intriguing open question.
@@ 1.1. Tautological classes on the moduli space of abelian varieties
$\lim_{g\to\infty}\mathsf{H}^{*}(\operatorname{Sp}_{2g}(\mathbb{Z}),\mathbb{Q})=\mathbb{Q}[\lambda_{1},\lambda_{3},\ldots]\,.$
@@ 1.3. Geometric cycles on $\mathcal{A}_{g}$
Formulas for $\mathsf{taut}([\mathcal{J}_{g}])$ for $g\leq 11$ can be found at [60].
@@ 1.4. The homomorphism property
By equation (3), if either $\alpha\in\mathsf{R}^{*}(\mathcal{A}_{g})$ or $\beta\in\mathsf{R}^{*}(\mathcal{A}_{g})$, then the homomorphism property holds for $(\alpha,\beta)$. The homomorphism property is therefore an obstruction to being tautological.
@@
    If $\beta$ is supported on the Noether-Lefschetz locus of abelian varieties that decompose (up to isogeny) as a product of $m$ elliptic curves and a (gm)(g-m)-dimensional abelian variety, then the homomorphism property holds for $(\alpha,\beta)$ for all $\alpha\in\mathsf{CH}^{*}(\mathcal{A}_{g})$, see [26]. In particular, the homomorphism property holds for $(\alpha,[\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{1}\times\mathcal{A}_{g-m}])$.
@@
We use two different approaches to compute (10). The first is a direct study of the excess intersection geometry of the fiber product of the Torelli and product morphisms following [10]. The second approach is by recasting the intersection geometry in terms of families Gromov-Witten classes, which are computed by a wall-crossing formula related to unramified maps [51, 49].
@@ 1.6. Wall-crossing
The wall-crossing formula established in [51, Theorem 6.21] relates the virtual fundamental classes of moduli spaces of stable and unramified maps. Let
@@
The partition $\mu=(1,1)$ arises for unramified maps to fibers of $\pi_{2}$ given by products $E_{1}\times E_{2}$ of elliptic curves: the associated domains curves are two disconnected genus 1 curves mapping to a nodal genus 2 curve consisting of two elliptic components. The point on the target corresponding to the node of the genus 2 curve has multiplicity $(1,1)$.
@@
$\bullet$ For $r\geq 3$, only partitions of the type $\mu=(1,\ldots,1)$ can occur since the minimal curves can have at worst ordinary multiple point singularities. The analysis for $r\geq 3$ is more complex and will not be treated here.
@@ 1.7. Results about the homomorphism property
The codimension of the product cycle $\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}$ in $\mathcal{A}_{g}$ is
@@
(15) $\mathsf{taut}([\mathcal{J}_{g}]\cdot[\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}])\stackrel{{\scriptstyle?}}{{=}}\mathsf{taut}([\mathcal{J}_{g}])\cdot\mathsf{taut}([\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}])\,\in\,\mathsf{R}^{\mathsf{cod}_{\mathcal{J}}+\mathsf{cod}_{\underline{g}}}(\mathcal{A}_{g})$
@@ 1.8. Conjectures
For $g\geq 3$, there are no such subvarieties $V\subset\mathcal{A}_{g}$ over the complex numbers [43]. However, in characteristic $p$, the locus $V_{0}\subset\mathcal{A}_{g}$ of abelian varieties with $p$-rank 0 provides a compact subvariety of the correct codimension (which represents a multiple of $\lambda_{g}$), see [52, 63]. Unfortunately, the Chow groups of $V_{0}$ have not yet been determined. Nevertheless, the existence of $V_{0}$ in characteristic $p$ may be viewed as motivation for the following weakening of Conjecture 8:
Conjecture 9 ($\overline{\mathbb{Q}}$ form).
The homomorphism property holds for all classes $\alpha,\beta\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ defined over $\overline{\mathbb{Q}}$.
Both Conjectures 8 and 9 are rather bold interpretations of the current data. Another possibility is that the homomorphism property holds in the restricted case where at least one of the classes $\alpha$ or $\beta$ is a Noether-Lefschetz cycle or, more generally, a special cycle  [^9] See [47] for a discussion of special cycles in the context of $\mathcal{A}_{g}$..
@@ 1.9. Stable curves and toroidal compactifications of $\mathcal{A}_{g}$
At the moment, Question 10 is open. In principle, both excess intersection theory and the wall-crossing approach can be applied. For the former, the scheme theoretic structure of the fiber product (17) over the divisor $\Delta_{0}\subset\overline{\mathcal{M}}_{g}$ would have to be determined. The wall-crossing approach has an advantage in that the families Gromov-Witten theory is well-defined for the family
@@ 1.10. The geometry of the universal fiber product $\mathcal{X}^{s}_{g}$
A difficulty in studying Question 10 and extending Theorem 11 to higher torus rank loci is the complexity of the boundary geometry of toroidal compactifications $\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}$. To generate elements of the Gorenstein kernel of $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g-s,2s})$ for general $s$, a different approach can be pursued using the $s$-fold fiber product
@@
The construction of $\operatorname{aj}^{*}(\Delta_{g,s})$ for $s>1$ may be viewed as probing the higher torus rank loci of toroidal compactifications. That $\operatorname{aj}^{*}(\Delta_{g,s})$ lies in the Gorenstein kernel was proven for $s=0$ in [10] and is proven for $s=1$ in Theorem 47, Section 6. The non-vanishing part of Conjecture 16 is open for all $s\geq 0$.
@@ 1.11. Gorenstein kernel
    $\pi_{i*}:\mathsf{K}_{g,n}\rightarrow\mathsf{K}_{g,n-1}$, where $\pi_{i}:\mathcal{M}_{g,n}^{\mathrm{ct}}\rightarrow\mathcal{M}_{g,n-1}^{\mathrm{ct}}$ is the map forgetting the $i^{th}$ marking,
@@
    $\pi^{*}_{i}:\mathsf{K}_{g,n-1}\rightarrow\mathsf{K}_{g,n}$,
@@
$\eta:\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1}+g_{2},n_{1}+n_{2}})\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})\,\,,\ \ \ \eta(\alpha\otimes\beta)=\pi_{1*}\left(\xi^{*}(\alpha)\cup\pi_{2}^{*}(\beta)\right)\,$,
where $\pi_{i}:\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1}\times\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{i},n_{i}+1}$ are the projections and
$\pi_{1*}:\mathsf{R}^{*}(\mathcal{M}_{g_{1},n_{1}+1}^{\mathrm{ct}})\times\mathsf{R}^{*}(\mathcal{M}_{g_{2},n_{2}+1}^{\mathrm{ct}})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})$
@@ 1.12. Acknowledgments
The calculation of $\mathsf{Tor}^{*}[\mathcal{A}_{2}\times\mathcal{A}_{g-2}]\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$ with the goal of determining whether the homomorphism property holds for $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ was a group project at ETH Zürich in the Fall term of 2025 with both theoretical and computational sides. We dedicate our paper to Carel Faber: his work [24] is the starting point of several of the themes developed here, and he pioneered the computational exploration of the tautological ring of the moduli space of curves.
@@ 4.4. Definition of $\psi$-classes
Up to certain normalization terms, these $\Psi$-classes can also be defined via vector bundles associated with the cotangent spaces at the markings, see [50, Section 3.2] for more details. The same definition applies to $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\circ}$.
@@ 5.2. The tautological ring of products of abelian varieties
$\varphi^{*}\theta=\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}\in\mathsf{CH}^{1}(X^{s})\,.\qed$
@@
By Proposition 33, it suffices to show that $\mathfrak{I}_{g,s}$ is Gorenstein. We decompose the cohomology $\mathsf{H}^{*}(X^{s},\mathbb{Q})$ into sums of irreducible representations of $\operatorname{Sp}(V),$ with $\mathfrak{I}_{g,s}$ occurring as the invariant summand. The top degree part $\mathsf{H}^{2gs}(X^{s})$ has rank $1$ and is spanned by the $\operatorname{Sp}(V)$-invariant class

### v2

@@ Abstract.
was constructed in [9]. We present here new calculations of intersection products of the Torelli locus in $\mathcal{A}_{g}$ with the product loci $\mathcal{A}_{r}\times\mathcal{A}_{g-r}\rightarrow\mathcal{A}_{g}$ for $r\leq 3$. The results suggest that $\mathsf{taut}$ is a $\mathbb{Q}$-algebra homomorphism (at least for special cycles). We discuss a conjectural framework for the homomorphism property of tautological projection.
@@
We define new tautological projections of cycles on the fiber products $\mathcal{X}_{g}^{s}\to\mathcal{A}_{g}$ of the universal family. We compute these projections for a class of product cycles on $\mathcal{X}_{g}^{s}$ in terms of a determinant involving the universal theta divisors and Poincaré classes. The Torelli pullback of the morphism $\mathcal{A}_{1}\times\mathcal{A}_{5}\to\mathcal{A}_{6}$ was used in [10] to construct a nontrivial element of the Gorenstein kernel of ${\mathsf{R}}^{5}(\mathcal{M}^{\mathrm{ct}}_{6})$. Using Abel-Jacobi pullbacks of product cycles on $\mathcal{X}_{g}^{s}$ and their projections, we construct a new family of classes which we conjecture to lie in the Gorenstein kernels of the tautological rings $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})$. In particular, we construct nontrivial elements of the Gorenstein kernels of $\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathrm{ct}})$ and $\mathsf{R}^{5}(\mathcal{M}_{4,4}^{\mathrm{ct}})$. Whether all of these classes generate the Gorenstein kernels of all $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ is an intriguing open question.
@@ 1.1. Tautological classes on the moduli space of abelian varieties
$\lim_{g\to\infty}\mathsf{H}^{*}(\operatorname{Sp}_{2g}(\mathbb{Z}),\mathbb{Q})=\mathbb{Q}[\lambda_{1},\lambda_{3},\lambda_{5},\ldots]\,.$
@@ 1.3. Geometric cycles on $\mathcal{A}_{g}$
Formulas for $\mathsf{taut}([\mathcal{J}_{g}])$ for $g\leq 11$ are available at the link [61].
@@ 1.4. The homomorphism property
By equation (3), if either $\alpha\in\mathsf{R}^{*}(\mathcal{A}_{g})$ or $\beta\in\mathsf{R}^{*}(\mathcal{A}_{g})$, then the homomorphism property holds for $(\alpha,\beta)$. Failure of the homomorphism property is therefore an obstruction to being tautological.
@@
    If $\beta$ is supported on the Noether-Lefschetz locus of abelian varieties that decompose (up to isogeny) as a product of $m>0$ elliptic curves and a (gm)(g-m)-dimensional abelian variety, then the homomorphism property holds for $(\alpha,\beta)$ for all $\alpha\in\mathsf{CH}^{*}(\mathcal{A}_{g})$, see [26]. In particular, the homomorphism property holds for $(\alpha,[\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{1}\times\mathcal{A}_{g-m}])$.
@@
We use two different approaches to compute (10). The first is a direct study of the excess intersection geometry of the fiber product of the Torelli and product morphisms following [10]. The second approach is by recasting the geometry in terms of families Gromov-Witten classes, which are computed by a wall-crossing formula related to unramified maps [52, 50].
@@ 1.6. Wall-crossing
The wall-crossing formula established in [52, Theorem 6.21] relates the virtual fundamental classes of moduli spaces of stable and unramified maps. Let
@@
The partition $\mu=(1,1)$ arises for unramified maps to fibers of $\pi_{2}$ given by products $E_{1}\times E_{2}$ of elliptic curves: the associated domain curves are two disconnected genus 1 curves mapping to a nodal genus 2 curve consisting of two elliptic components. The point on the target corresponding to the node of the genus 2 curve has multiplicity $(1,1)$.
Since curves in the minimal class $\beta_{\mathrm{min}}$ are at worst nodal, maps to an abelian surface can only have multiplicities $(1)$ and $(1,1)$. By admissibility, the curves have degree at most $2$ over rational components of Fulton-MacPherson degenerations of abelian surfaces. Degree $2$ curves in $\mathbb{P}^{2}$ are also at worst nodal. Hence, the same multiplicities $(1)$ and $(1,1)$ occur for maps to Fulton-MacPherson degenerations.
@@ 1.7. Results about the homomorphism property
The codimension of the product cycle $\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}$ in $\mathcal{A}_{g}$ is
@@
(15) $\mathsf{taut}([\mathcal{J}_{g}]\cdot[\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\stackrel{{\scriptstyle?}}{{=}}\mathsf{taut}([\mathcal{J}_{g}])\cdot\mathsf{taut}([\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\,\in\,\mathsf{R}^{\mathsf{cod}_{\mathcal{J}}+\mathsf{cod}_{\underline{g}}}(\mathcal{A}_{g})$
@@ 1.8. Conjectures
For $g\geq 3$, there are no such subvarieties $V\subset\mathcal{A}_{g}$ over the complex numbers [43]. However, in characteristic $p$, the locus $V_{0}\subset\mathcal{A}_{g}$ of abelian varieties with $p$-rank 0 provides a compact subvariety of the correct codimension (which represents a multiple of $\lambda_{g}$), see [53, 64]. Unfortunately, the Chow groups of $V_{0}$ have not yet been determined. A specialization argument shows that the homomorphism property for cycles on $\mathcal{A}_{g}(\overline{\mathbb{F}}_{p})$ for all $p$ implies the homomorphism property for cycles on $\mathcal{A}_{g}(\overline{\mathbb{Q}})$. Since every cycle over $\mathbb{C}$ is homologous to a cycle defined over $\overline{\mathbb{Q}}$, Conjecture 8 would then follow.
Conjecture 8 is a rather bold interpretation of the current data. Another possibility is that the homomorphism property holds in the restricted case where at least one of the classes $\alpha$ or $\beta$ is a Noether-Lefschetz cycle or, more generally, a special cycle  [^9] See [47] for a discussion of special cycles in the context of $\mathcal{A}_{g}$..
@@ 1.9. Stable curves and toroidal compactifications of $\mathcal{A}_{g}$
At the moment, Question 9 is open. In principle, both excess intersection theory and the wall-crossing approach can be applied. For the former, the scheme theoretic structure of the fiber product (17) over the divisor $\Delta_{0}\subset\overline{\mathcal{M}}_{g}$ would have to be determined. The wall-crossing approach has an advantage in that the families Gromov-Witten theory is well-defined for the family
@@ 1.10. The geometry of the universal fiber product $\mathcal{X}^{s}_{g}$
A difficulty in studying Question 9 and extending Theorem 10 to higher torus rank loci is the complexity of the boundary geometry of toroidal compactifications $\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}$. To generate elements of the Gorenstein kernel of $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g-s,2s})$ for general $s$, a different approach can be pursued using the $s$-fold fiber product
@@
The construction of $\operatorname{aj}^{*}(\Delta_{g,s})$ for $s>1$ may be viewed as probing the higher torus rank loci of toroidal compactifications. That $\operatorname{aj}^{*}(\Delta_{g,s})$ lies in the Gorenstein kernel was proven for $s=0$ in [10] and is proven for $s=1$ in Theorem 48 of Section 6. The class $\operatorname{aj}^{*}(\Delta_{4,2})\in\mathsf{R}^{5}(\mathcal{M}_{4,4}^{\mathsf{ct}})$ is proven to be a nontrivial element of the kernel of the $\lambda_{4}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{4,4}^{\mathrm{ct}})$ in Theorem 48. The non-vanishing part of Conjecture 15 is open for all $s\geq 0$.
@@ 1.11. Gorenstein kernel
    $\rho_{i*}:\mathsf{K}_{g,n}\rightarrow\mathsf{K}_{g,n-1}$, where $\rho_{i}:\mathcal{M}_{g,n}^{\mathrm{ct}}\rightarrow\mathcal{M}_{g,n-1}^{\mathrm{ct}}$ is the map forgetting the $i^{th}$ marking,
@@
    $\rho^{*}_{i}:\mathsf{K}_{g,n-1}\rightarrow\mathsf{K}_{g,n}$,
@@
$\eta:\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1}+g_{2},n_{1}+n_{2}})\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})\,\,,\ \ \ \eta(\alpha\otimes\beta)=\rho_{1*}\left(\xi^{*}(\alpha)\cup\rho_{2}^{*}(\beta)\right)\,$,
where $\rho_{i}:\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1}\times\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{i},n_{i}+1}$ are the projections and
$\rho_{1*}:\mathsf{R}^{*}(\mathcal{M}_{g_{1},n_{1}+1}^{\mathrm{ct}})\times\mathsf{R}^{*}(\mathcal{M}_{g_{2},n_{2}+1}^{\mathrm{ct}})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})$
@@ 1.12. Acknowledgments
The calculation of $\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$ with the goal of determining whether the homomorphism property holds for $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ was a group project at ETH Zürich in the Fall term of 2025 with both theoretical and computational sides. We dedicate our paper to Carel Faber: his work [24] is the starting point of several of the themes developed here, and he pioneered the computational exploration of the tautological ring of the moduli space of curves.
@@ 4.4. Definition of $\psi$-classes
Up to certain normalization terms, these $\Psi$-classes can also be defined via vector bundles associated with the cotangent spaces at the markings, see [51, Section 3.2] for more details. The same definition applies to $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\circ}$.
@@ 5.2. The tautological ring of products of abelian varieties
(42) $\varphi^{*}\theta=\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}\in\mathsf{CH}^{1}(X^{s})\,.\qed$
@@
By Proposition 32, it suffices to show that $\mathfrak{I}_{g,s}$ is Gorenstein. Since $\operatorname{Sp}(V)$ is reductive, we can decompose the cohomology $\mathsf{H}^{*}(X^{s},\mathbb{Q})$ into sums of irreducible representations, with $\mathfrak{I}_{g,s}$ occurring as the invariant summand. The top degree part $\mathsf{H}^{2gs}(X^{s})$ has rank $1$ and is spanned by the $\operatorname{Sp}(V)$-invariant class
@@ 5.3. Tautological projection
5.4. Properties of tautological rings
By definition, tautological rings are preserved under pullbacks via projections forgetting a factor:
(43) $\text{pr}:X^{n+1}\to X^{n}\,,\ \ \ \ \text{pr}^{*}:\mathsf{R}^{*}(X^{n})\rightarrow\mathsf{R}^{*}(X^{n+1})\,.$
By [3, Theorem 9.10], tautological rings are also preserved under the corresponding pushforwards,
(44) $\text{pr}_{*}:\mathsf{R}^{*}(X^{n+1})\rightarrow\mathsf{R}^{*}(X^{n})\,.$
In fact, much more general versions of the properties (43) and (44) hold.
Let $\mathsf{A}=(a_{ij})$ be a matrix of size $t\times s$ with integer coefficients. The matrix $\mathsf{A}$ induces a natural morphism
(45) $f_{\mathsf{A}}:X^{s}\to X^{t},\quad(x_{1},\ldots,x_{s})\to\left(\sum_{j=1}^{s}a_{1j}x_{j},\ldots,\sum_{j=1}^{s}a_{tj}x_{j}\right)\,.$
The diagonal inclusion $X\rightarrow X\times X$ is obtained from the case $\mathsf{A}=\begin{pmatrix}1\\ 1\end{pmatrix}\,$.
Proposition 36.
The tautological rings are preserved under pushforward,
$(f_{\mathsf{A}})_{*}:\mathsf{R}^{*}(X^{s})\rightarrow\mathsf{R}^{*}(X^{t})\,.$
Proof.
For all classes $\alpha\in\mathsf{CH}^{*}(X^{s})$, we have
$(f_{\mathsf{A}})_{*}\alpha=\widehat{\text{pr}}_{*}\left({\text{pr}}^{*}\alpha\cdot\left[\Gamma_{\mathsf{A}}\right]\right)\,$,
where $\Gamma_{\mathsf{A}}\subset X^{s+t}$ is the graph of $f_{\mathsf{A}}$, and $\text{pr},$ $\widehat{\text{pr}}$ are the projections to $X^{s}$ and $X^{t}$ respectively. By the pushforward property (44), if $\left[\Gamma_{\mathsf{A}}\right]\in\mathsf{R}^{*}(X^{s+t})$, then $(f_{\mathsf{A}})_{*}\alpha\in\mathsf{R}^{*}(X^{t})$. We may further assume $t=1$ by using the following equation:
$\left[\Gamma_{\mathsf{A}}\right]=\widehat{\text{pr}}_{1}^{*}\left[\Gamma_{1}\right]\cdots\widehat{\text{pr}}_{t}^{*}\left[\Gamma_{t}\right]\,$,
where $\Gamma_{i}\subset X^{s+1}$ is the graph of the morphism
$(x_{1},\ldots,x_{s})\to\sum_{j=1}^{s}a_{ij}x_{j}\,.$
Here, for $1\leq i\leq t$, the projection $\widehat{\text{pr}}_{i}:X^{s+t}\to X^{s+1}$ records the first $s$ coordinates as well as the coordinate in position $s+i$.
For $t=1$, we write $(a_{1},\ldots,a_{s})$ for the entries of $\mathsf{A}$, and we will prove $\left[\Gamma_{\mathsf{A}}\right]\in\mathsf{R}^{g}(X^{s+1})$. Setting
$\varphi:X^{s+1}\to X,\quad(x_{1},\ldots,x_{s},x_{s+1})\to x_{s+1}-a_{1}x_{1}-\ldots-a_{s}x_{s}\,$,
we have
$\left[\Gamma_{\mathsf{A}}\right]=\varphi^{*}([0])=\varphi^{*}\left(\frac{\theta^{g}}{g!}\right)\,.$
Finally, to show $\varphi^{*}\theta\in\mathsf{CH}^{1}(X^{s+1})$ is tautological, we use the identity
$\varphi^{*}\theta=\theta_{s+1}+\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}-\sum_{i=1}^{s}a_{i}\eta_{i,s+1}\,$,
which is a particular case of (42). ∎
Proposition 37.
The tautological rings are preserved under pullback,
$(f_{\mathsf{A}})^{*}:\mathsf{R}^{*}(X^{t})\rightarrow\mathsf{R}^{*}(X^{s})\,.$
Proof.
By the definition of $\mathsf{R}^{*}(X^{t})$, we can reduce to the case where $t=1$ or $t=2$. The $t=1$ case is a consequence of (42). For $t=2$, we write
$\mathsf{A}=\begin{pmatrix}a_{1}&\ldots&a_{s}\\ b_{1}&\ldots&b_{s}\end{pmatrix}\,$,
and we use (42) and the equation
$(f_{\mathsf{A}})^{*}(\eta_{12})=\sum_{1\leq i<j\leq s}(a_{i}b_{j}+a_{j}b_{i})\,\eta_{ij}+2\sum_{i=1}^{s}a_{i}b_{i}\,\theta_{i}\,.$
The latter is easily verified by induction on $s$ and the See-Saw Theorem. ∎